Direct Labor Mix and Yield Variances

Delano Company uses two types of direct labor for the manufacturing of its products: fabricating and assembly. Delano has developed the following standard mix for direct labor, where output is measured in number of circuit boards.

Direct Labor Type Mix SP Standard
Cost
Fabricating 2 hrs. $25       $50   
Assembly 3 11       33   
Total 5 hrs.    $83     
Yield 25 units

During the second week in April, Delano produced the following results:

Direct Labor Type Actual Mix
Fabricating 25,000 hrs.
Assembly 50,000
Total 75,000 hrs.
Yield 370,000 units

Required:

Instructions for parts 3 and 4: If a variance is zero, enter "0" and select "Not applicable" from the drop down box.

1. Calculate the yield ratio.

2. Calculate the standard cost per unit of the yield. Round your answers to the nearest cent.
$ per unit of yield

3. Calculate the direct labor yield variance.
$  

4. Calculate the direct labor mix variance.
$  
Answer #1
yield ratio
TOTAL UNITS $        370,000
TOTAL LABOR HOURS        75,000.00
                  4.93
Direct Labor Type Mix SP Standard
Cost
Fabricating 2 hrs. $25       $50
Assembly 3 11       $33
Total 5 hrs.    $83
Yield 25 units
COST PER UNIT OF YIELD $       3.32
$3.00
DIRECT LABOR YIELD VARIANCE = (Actual units of labor inputs used – allowed for actual outputs) x (Standard average price per unit of labor input).
FABRICATING
Actual units of labor inputs used 25000 hrs
Allowed for actual outputs = 2/25*370000 29600 hrs
Standard average price per unit of labor input $25
DIRECT LABOR YIELD VARIANCE = (25000-29600) *25
$        (115,000) unfavorable
Assembly
Actual units of labor inputs used 50000 hrs
Allowed for actual outputs = 3/25*370000 44400 hrs
Standard average price per unit of labor input $11
DIRECT LABOR YIELD VARIANCE = (50000-44400)*11
$             61,600 favorable
Labor Mix Variance = (Revised Standard – Actual Hours Worked) X Standard Rate
Revised standard Hours (RSH)= Actual Total Hours Worked X Standard Ratio of Workers
FABRICATING
Actual units of labor inputs used 25000 hrs
Standard average price per unit of labor input $25
RSH = Actual Total Hours Worked X Standard Ratio of Workers = 75000*2/5 30000 hrs
Labour Mix Variance (30000-25000) * 25
$           125,000 FAVORABLE
Assembly
Actual units of labor inputs used 50000 hrs
Standard average price per unit of labor input $11
RSH = Actual Total Hours Worked X Standard Ratio of Workers = 75000*3/5 45000 hrs
Labour Mix Variance (45000-50000) * 11
55000 unfavorable